Circular No. 8026

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002jj IN IC 340 D. Hutchings and W. Li report the LOTOSS (cf. IAUC 7906) discovery

of an apparent supernova (mag ∼ 16.5) on an unfiltered KAIT image taken on Dec. 3.4 UT. The new object was confirmed at about the same brightness on images taken on Oct. 24.4 and Nov. 3.4. SN 2002jj is located at α = 3^{h}39^{m}29^{s}.55, δ = −13^{o}06′55′′.8 (equinox 2000.0), which is 7′′.0 east and 1′′.5 south of the center of IC 340. Images taken on Oct. 1.4 (limiting mag ∼ 18) and Aug. 14.5 (limiting mag ∼ 19) showed nothing at this position.

SUPERNOVA 2002jk Further to IAUC 7902, W. M. Wood-Vasey et al. report the discovery of

an apparent supernova (mag 17.9) on an unfiltered NEAT image obtained on Feb. 14.37 UT with the Palomar 1.2-m Schmidt telescope. The new object was located at α = 10^{h}44^{m}14^{s}.39, δ = +15^{o}58′09′′.9 (equinox 2000.0), which is 1′′ east and 5′′ south of the center of the apparent host galaxy. Additional magnitudes for SN 2002jk: Jan. 21 and 22, [20.5 (S/N = 3); Feb. 4, 20.4: (possible detection); 14, 17.9; 23, 17.6; Mar. 11, 18.0; 22, 18.3; Apr. 22, 18.5.

SUPERNOVA 2002ji IN NGC 3655 A. Clocchiatti, Pontificia Universidad Católica de Chile, reports that

a fully reduced spectrogram (range 390–705 nm) of SN 2002ji (cf. IAUC 8025), obtained on Dec. 3.3 UT with the Baade 6.5-m telescope (+ Boller & Chivens spectrograph) at Las Campanas Observatory, shows it to be of type Ic. A comparison with spectra of SNe 1994I and 1983V indicates that SN 2002ji is 3–6 weeks past maximum light.

COMETS C/1999 O4; C/2002 U6, U7, U8, U9 (SOHO) Further to IAUC 8025, D. Hammer reports measurements for a comet

near the sun found by R. Kracht on SOHO website C3 coronagraph images. The reduced astrometry (discovery position below) and orbital elements (showing no known group identity) by B. G. Marsden appear on MPEC 2002-W34. Regarding the new SOHO comets on IAUC 8025, C/2002 U6 belongs to the Meyer group, while the rest are Kreutz sungrazers.

1999 UT  α_{2000}  δ_{2000}
July 18.821  7^{h}44.^{m}7  +23^{o}40′
2002 December 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green